# Stability of a passive viscous droplet in a confined active nematic liquid crystal

### v1

@@ I Introduction
The interaction of active materials with passive phases is relevant to many biological systems spanning a range of length scales. Examples on large scales include: front propagation in spreading bacterial swarms and suspensions, where complex interfacial instabilities have been observed Patteson et al. (2018); Martínez-Calvo et al. (2022); cell migration during the formation of cancer metastasis Nürnberg et al. (2011); Hoshino et al. (2013); or the migration of confluent epithelial cell layers during wound healing Martin and Parkhurst (2004); Armengol-Collado et al. (2023). On cellular length scales, the organization of various subcellular compartments and organelles hinges on their interaction with the active cytoskeleton and cytoplasmic flows Chu et al. (2017); Mogilner and Manhart (2018). On yet smaller scales, the placement and structural organization of subnuclear bodies and heterochromatin compartments inside the cell nucleus have been hypothesized to be affected by the dynamics of transcriptionally active euchromatin Caragine et al. (2018); Mahajan et al. (2022). Interfaces between active and passive phases have also been of interest in synthetic and reconstituted systems, e.g., in active nematic fluids composed of microtubules and molecular motors Sanchez et al. (2012); Shelley (2016), where activity has been shown to destabilize interfaces and drive shape fluctuations and active waves Adkins et al. (2022).
@@
In all of these systems, the dynamics and emerging morphologies of active/passive interfaces stem from the interplay of active, elastic, viscous, and possibly capillary stresses. A perturbation to the interfacial shape affects the orientation of the active matter constituents, thus giving rise to both active and elastic stresses. These stresses in turn result in tractions on the interface as well as flows in the material, whose net effect can be either stabilizing or destabilizing. A few theoretical and computational models have been proposed in recent years to explain these effects. The dynamics of active droplets, either in free space or near boundaries, has been studied numerically and analytically, where active stresses and the flows they induce can lead to spontaneous propulsion, fingering instabilities, division or spreading Tjhung et al. (2012); Joanny and Ramaswamy (2012); Giomi and DeSimone (2014); Whitfield et al. (2014); Khoromskaia and Alexander (2015); Young et al. (2021); Alert (2022). Active films have also been considered, where interfacial instabilities can either be triggered or stabilized by activity depending on the sign of active stresses Voituriez et al. (2006); Sankararaman and Ramaswamy (2009); Blow et al. (2017); Alonso-Matilla and Saintillan (2019). Other studies have analyzed the motion of passive objects, either rigid or deformable, suspended in active suspensions, where active stresses can give rise to spontaneous transport Freund (2023, 2024), rotation Fürthauer et al. (2012), and deformations Chandler and Spagnolie (2024).
@@
The paper is organized as follows. We formulate the problem in Sec. II, where we present the governing equations, boundary conditions, and non-dimensionalization. Section III presents details of the linear stability analysis of the axisymmetric base state and a discussion of the dispersion relation and growth rates. The inverted problem, i.e., the stability of an active nematic drop within a passive layer and its associated dispersion relation are discussed in Sec. IV. We conclude in Sec. V.
II Problem formulation
@@
We analyze the stability of a viscous Newtonian droplet surrounded by a sharply aligned active nematic fluid in circular confinement in two dimensions (Fig. 1). The system is enclosed by a circular boundary of radius $R$ centered at the origin $O$. In polar coordinates $(r,\theta)$, the shape of the interface between the drop and nematic layer is described by the Monge representation $r=\eta R+\zeta(\theta,t)$, where $\eta\in(0,1)$ is the dimensionless radius ratio, and $\zeta$ captures deviations from the reference base state where the drop is circular and centered at the origin. We denote by $\Omega^{-}$ the domain occupied by the passive drop, by $\Omega^{+}$ the domain occupied by the nematic, and by $\mathcal{C}$ the curve describing the shape of the interface between them.
@@ II.1 Nematodynamics
Orientational order in the active nematic layer surrounding the droplet is described by a nematic director field $\mathbf{q}(\mathbf{x},t)$ (with $|\mathbf{q}|=1$ in the sharply aligned limit), which satisfies the Leslie-Ericksen model Ericksen (1961, 1962); Leslie (1968); De Gennes and Prost (1995); Gao et al. (2017)
@@
where $\mathbf{u}$ is the fluid velocity in the nematic. The rate-of-strain tensor $\mathbf{e}=(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})/2$ and rate-of-rotation tensor $\mathbf{w}=(\nabla\mathbf{u}-\nabla\mathbf{u}^{T})/2$ depend on the velocity gradient defined as $(\nabla\mathbf{u})_{ij}=\partial_{i}u_{j}$, and $\varpi\in[0,1]$ denotes the flow alignment parameter. The last term in Eq. (1) captures nematic elasticity, and involves the rotational viscosity $\Gamma$ and molecular field $\mathbf{h}$, which derives from a free energy density. Under the commonly used one-constant approximation De Gennes and Prost (1995), we express the free energy density $\mathcal{F}_{d}$ as
@@
Nematic elasticity thus enters Eq. (1) as a diffusive term, with the ratio $K/\Gamma$ playing the role of a diffusivity. The nematic is assumed to satisfy strong parallel anchoring conditions at both the outer domain boundary and droplet interface. Note that in two dimensions the director field is equivalently described by a single angle $\beta(\mathbf{x},t)$ such that $\mathbf{q}=[\cos\beta,\sin\beta]$; we make further use of this representation below.
@@ II.2 Mass and momentum balance
$\nabla\cdot\mathbf{u}=0,\qquad\nabla\cdot(\bm{\sigma}^{h}+\bm{\sigma}^{e}+\bm{\sigma}^{a})=\bm{0},$ (4)
where $\bm{\sigma}^{h}$, $\bm{\sigma}^{a}$ and $\bm{\sigma}^{e}$ are the hydrodynamic, active, and elastic (or Leslie-Ericksen) stress tensors, respectively:
@@
$\bm{\sigma}^{e}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\smash{\left(\tfrac{1}{2}K\|\nabla\mathbf{q}\|^{2}-\varpi\Gamma\mathbf{qq}\,:\,\mathbf{e}\right)\mathbf{I}-K\nabla\mathbf{q}^{T}\cdot\nabla\mathbf{q}-K\left[(\nabla^{2}\mathbf{q})\mathbf{q}+\|\nabla\mathbf{q}\|^{2}\mathbf{qq}\right]+\varpi\Gamma(\mathbf{qq}\,:\,\mathbf{e})\mathbf{qq}.}}$ (7)
Here, $p$ is the fluid pressure and $\mu$ is the shear viscosity of the fluid. The active stress magnitude is $\alpha=n\varsigma_{a}$, where $n$ is the number density of active dipoles and $\varsigma_{a}$ is the dipole strength or stresslet; $\alpha>0$ for an extensile nematic, and $\alpha<0$ in the contractile case Saintillan and Shelley (2013); Saintillan (2018).
@@ II.3 Boundary conditions
To parametrize the shape of the drop, we introduce the function $F(r,\theta,t)=r-\eta R-\zeta(\theta,t)$ and note that the interface $\mathcal{C}$ between the two fluids is defined as $F=0$. Consequently, the outward unit normal to the interface is
@@
Note that the fluid velocity is continuous at the interface: $\mathbf{u}=\overline{\mathbf{u}}$ at $F=0$.
@@
$\mathbf{n}\cdot(\bm{\sigma}^{h}+\bm{\sigma}^{a}+\bm{\sigma}^{e}-\overline{\bm{\sigma}}^{h})=\gamma\kappa\mathbf{n},$ (12)
where $\gamma$ is the surface tension, assumed to be uniform, and $\kappa=\nabla\cdot\mathbf{n}$ is the local curvature. The parallel anchoring condition for the nematic at the drop surface reads: $\mathbf{q}\cdot\mathbf{n}=0$ at $F=0$.
@@ II.4 Non-dimensionalization
$\bm{\sigma}^{h}=-p\,\mathbf{I}+2\,\mathbf{e},\qquad\bm{\sigma}^{a}=-S\mathbf{qq},\qquad\bm{\sigma}^{e}=-\xi^{2}\nabla\mathbf{q}^{\mathrm{T}}\cdot\nabla\mathbf{q},$ (15)
@@
where $\lambda=\overline{\mu}/\mu$ is the viscosity ratio. The active stress coefficient in Eq. (15) is $S=\alpha/|\alpha|$, with $S=+1$ for an extensile nematic and $S=-1$ for a contractile nematic. The relative magnitude of elastic over active stresses is captured by the dimensionless parameter $\xi=\sqrt{K/|\alpha|}/R$, or ratio of the active length scale $\sqrt{K/|\alpha|}$ to the system size. Active stresses are negligible when $\xi\gg 1$, whereas they dominate when $\xi\ll 1$, a regime where active turbulence is typically observed. In the present work, we assume $\xi\sim O(1)$.
@@
$\mathbf{n}\cdot(\bm{\sigma}^{h}+\bm{\sigma}^{a}+\bm{\sigma}^{e}-\overline{\bm{\sigma}}^{h})=\frac{1}{\mathrm{Ca}}\kappa\mathbf{n},$ (17)
where we have introduced the active capillary number $\mathrm{Ca}=|\alpha|R/\gamma$ comparing the magnitudes of active vs. capillary stresses. Under the assumption of vanishing Péclet number, the system is entirely governed by five parameters: the radius ratio $\eta$, the sign $S$ of active stresses, the capillary number $\mathrm{Ca}$, the dimensionless elastic constant $\xi$, and the viscosity ratio $\lambda$.
@@ III Linear stability analysis
We analyze the stability of small perturbations from the stable case of a circular drop ($\zeta_{0}=0$) centered at the origin. In the following, subscripts 0 and 1 are used to denote the base state and linear perturbations, respectively.
@@ III.1 Circular base state
$\beta_{0}=\arctan\frac{y}{x}+\frac{\pi}{2}=\theta+\frac{\pi}{2},$ (18)
which satisfies Eq. (14) along with the anchoring boundary conditions at $r=\eta$ and $1$. The corresponding nematic director is $\mathbf{q}_{0}=\hat{\bm{\theta}}$. Both fluids are still ($\mathbf{u}_{0}=\overline{\mathbf{u}}_{0}=\mathbf{0}$). The base-state active and elastic stress tensors in the outer nematic layer are given by
$\bm{\sigma}^{a}_{0}=-S\mathbf{q}_{0}\mathbf{q}_{0}=-S\hat{\bm{\theta}}\hat{\bm{\theta}},\qquad\bm{\sigma}^{e}_{0}=-\xi^{2}\nabla\mathbf{q}_{0}^{\mathrm{T}}\cdot\nabla\mathbf{q}_{0}=-\frac{\xi^{2}}{r^{2}}\hat{\mathbf{r}}\hat{\mathbf{r}}.$ (19)
@@
$\overline{p}_{0}=S\ln\eta+\frac{\xi^{2}}{2\eta^{2}}+\frac{1}{\eta\mathrm{Ca}}.$ (22)

### v2

# Stability of co-annular active and passive confined fluids

@@ I Introduction
The interaction of active materials with passive phases is relevant to many biological systems spanning a range of length scales. Examples on large scales include: front propagation in spreading bacterial swarms and suspensions, where complex interfacial instabilities have been observed Patteson et al. (2018); Mart´ınez-Calvo et al. (2022); cell migration during the formation of cancer metastasis Nürnberg et al. (2011); Hoshino et al. (2013); or the migration of confluent epithelial cell layers during wound healing Martin and Parkhurst (2004); Armengol-Collado et al. (2023). On cellular length scales, the organization of various subcellular compartments and organelles hinges on their interaction with the active cytoskeleton and cytoplasmic flows Chu et al. (2017); Mogilner and Manhart (2018). On yet smaller scales, the placement and structural organization of subnuclear bodies and heterochromatin compartments inside the cell nucleus have been hypothesized to be affected by the dynamics of transcriptionally active euchromatin Caragine et al. (2018); Mahajan et al. (2022); Rautu et al. (2025). Interfaces between active and passive phases have also been of interest in synthetic and reconstituted systems, e.g., in active nematic fluids composed of microtubules and molecular motors Sanchez et al. (2012); Shelley (2016), where activity has been shown to destabilize interfaces and drive shape fluctuations and active waves Adkins et al. (2022).
@@
In all of these systems, the dynamics and emerging morphologies of active/passive interfaces stem from the interplay of active, elastic, viscous, and capillary stresses. A perturbation to the interfacial shape affects the orientation of the active matter constituents, thus giving rise to both active and elastic stresses. These stresses in turn result in tractions on the interface as well as flows in the material, whose net effect can be either stabilizing or destabilizing. A few theoretical and computational models have been proposed in recent years to explain these effects. The dynamics of active droplets, either in free space or near boundaries, has been studied numerically and analytically, where active stresses and the flows they induce can lead to spontaneous propulsion, fingering instabilities, division or spreading Tjhung et al. (2012); Joanny and Ramaswamy (2012); Giomi and DeSimone (2014); Whitfield et al. (2014); Khoromskaia and Alexander (2015); Young et al. (2021); Alert (2022). Active films have also been considered, where interfacial instabilities can either be triggered or stabilized by activity depending on the sign of active stresses Voituriez et al. (2006); Sankararaman and Ramaswamy (2009); Blow et al. (2017); Alonso-Matilla and Saintillan (2019). Other studies have analyzed the motion of passive objects, either rigid or deformable, suspended in active suspensions, where active stresses can give rise to spontaneous transport Freund (2023, 2024), rotation Fürthauer et al. (2012), and deformations Chandler and Spagnolie (2024a).
@@
The paper is organized as follows. We formulate the problem in Sec. II, where we present the governing equations, boundary conditions, and non-dimensionalization. Section III presents details of the linear stability analysis of the axisymmetric base state and a discussion of the dispersion relation and growth rates. The inverted problem, i.e., the stability of an active nematic drop within a passive layer and its associated dispersion relation, is discussed in Sec. IV. We conclude in Sec. V.
II System formulation
@@
We analyze the stability of a viscous Newtonian droplet surrounded by a sharply aligned active nematic fluid in circular confinement in two dimensions (Fig. 1). The system is enclosed by a circular boundary of radius $R$ centered at the origin $O$. In polar coordinates $(r,\theta)$ with unit basis vectors $\smash{(\hat{\mathbf{r}},\hat{\bm{\theta}})}$, the shape of the interface between the drop and nematic layer is described by the Monge representation $r=\eta R+\zeta(\theta,t)$, where $\eta\in(0,1)$ is the dimensionless radius ratio, and $\zeta$ captures deviations from the reference base state where the drop is circular and centered at the origin. We denote by $\Omega^{-}$ the domain occupied by the passive drop, by $\Omega^{+}$ the domain occupied by the nematic, and by $\mathcal{C}$ the curve describing the shape of the interface between them.
@@ II.1 Nematodynamics
Orientational order in the active nematic layer surrounding the droplet is described by a nematic director field $\mathbf{q}(\mathbf{x},t)$ (with $|\mathbf{q}|=1$ in the sharply aligned limit), which satisfies the Ericksen–Leslie model Ericksen (1961, 1962); Leslie (1968); Laudau et al. (1986); De Gennes and Prost (1995); Stewart (2004); Gao et al. (2017)
@@
where $\mathbf{u}$ is the fluid velocity in the nematic. The rate-of-strain tensor $\mathbf{e}=(\nabla\mathbf{u}+\nabla\mathbf{u}^{\mathrm{T}})/2$ and rate-of-rotation tensor $\mathbf{w}=(\nabla\mathbf{u}-\nabla\mathbf{u}^{\mathrm{T}})/2$ depend on the velocity gradient defined as $(\nabla\mathbf{u})_{ij}=\partial_{i}u_{j}$, and $\varpi$ denotes the flow alignment parameter. The last term in Eq. (1) captures nematic elasticity, and involves the rotational viscosity $\Gamma$ and molecular field $\mathbf{h}$, which derives from a free energy density. Under the commonly used one-constant approximation De Gennes and Prost (1995), we express the free energy density $\mathcal{F}_{d}$ as
@@
Nematic elasticity thus enters Eq. (1) as a diffusive term, with the ratio $K/\Gamma$ playing the role of a diffusivity. The nematic is assumed to satisfy strong parallel anchoring conditions at both the outer domain boundary and droplet interface. In two dimensions, the director field is equivalently described by a single angle $\beta(\mathbf{x},t)$ such that $\mathbf{q}=[\cos\beta,\sin\beta]$; we make further use of this representation below.
@@ II.2 Mass and momentum balance
$\nabla\cdot\mathbf{u}=0,\qquad\nabla\cdot(\bm{\sigma}^{h}+\bm{\sigma}^{a}+\bm{\sigma}^{e})=\bm{0},$ (4)
where $\bm{\sigma}^{h}$, $\bm{\sigma}^{a}$ and $\bm{\sigma}^{e}$ are the hydrodynamic, active, and elastic (or Ericksen–Leslie) stress tensors, respectively:
@@
$\begin{split}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\sigma}^{e}}&\,\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\big{(}\tfrac{1}{2}K\|\nabla\mathbf{q}\|^{2}-\varpi\Gamma\mathbf{pp}:\mathbf{e}\big{)}\mathbf{I}-K\nabla\mathbf{q}\cdot\nabla\mathbf{q}^{\mathrm{T}}+\varpi^{2}\Gamma(\mathbf{qq}\,:\,\mathbf{e})\mathbf{qq}}\\ &\,\,\,\,\,\,\,\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\tfrac{1}{2}K\left[(\varpi+1)\mathbf{q}(\nabla^{2}\mathbf{q})+(\varpi-1)(\nabla^{2}\mathbf{q})\mathbf{q}+2\varpi\|\nabla\mathbf{q}\|^{2}\mathbf{qq}\right].}\end{split}$ (7)
Here, $p$ is the fluid pressure, and $\mu$ is the shear viscosity of the fluid. The active stress magnitude is $\alpha=n\varsigma_{a}$, where $n$ is the number density of active dipoles and $\varsigma_{a}$ is the dipole strength or stresslet; $\alpha>0$ for an extensile nematic, and $\alpha<0$ in the contractile case Saintillan and Shelley (2013); Saintillan (2018). We note that several expressions for $\bm{\sigma}^{e}$ with varying levels of complexity exist in the literature Laudau et al. (1986); De Gennes and Prost (1995); Stewart (2004). The version provided in Eq. (7) assumes an isotropic viscous response; see Cates & Tjhung Cates and Tjhung (2018) for a derivation. In Eq. (4), we adopt the convention $(\nabla\cdot\bm{\sigma})_{i}=\partial_{k}\sigma_{ki}$ for the tensorial divergence; accordingly, the traction on a surface with outward unit normal $\mathbf{n}$ is defined as $\mathbf{t}=\mathbf{n}\cdot\bm{\sigma}$.
@@ II.3 Boundary conditions
To parametrize the shape of the drop, we introduce the function $F(r,\theta,t)=r-\eta R-\zeta(\theta,t)$ so that the interface $\mathcal{C}$ between the two fluids is defined as $F=0$. Consequently, the outward unit normal to the interface is
@@
Finally, the fluid velocity is continuous at the interface: $\mathbf{u}=\overline{\mathbf{u}}$ at $F=0$.
@@
$\mathbf{n}\cdot(\bm{\sigma}^{h}+\bm{\sigma}^{a}+\bm{\sigma}^{e}-\overline{\bm{\sigma}}^{h})+\mathbf{t}^{e}_{s}=\gamma\kappa\mathbf{n},$ (12)
where $\gamma$ is the uniform surface tension, and $\kappa=\nabla\cdot\mathbf{n}$ is the local curvature. The parallel anchoring condition for the nematic at the drop surface reads $\mathbf{q}\cdot\mathbf{n}=0$ at $F=0$, and results in an additional elastic surface traction $\mathbf{t}^{e}_{s}$ in Eq. (12) expressed as Chandler and Spagnolie (2024b):
$\mathbf{t}^{e}_{s}=K\frac{\partial}{\partial s}\bigg{(}\frac{\partial\beta}{\partial n}\mathbf{n}\bigg{)}.$ (13)
Here, $\partial/\partial n$ and $\partial/\partial s$ denote the normal and arc-length derivatives along the interface, respectively, with the normal pointing into the nematic and the direction of arclength determined by a ninety-degree counterclockwise rotation of the normal.
The choice of strong anchoring boundary conditions at the domain boundaries is convenient mathematically as it simplifies the solution of the boundary problem for $\beta$. The choice of parallel versus normal anchoring, however, is the most physically relevant as many of the systems that motivate this work (such as microtubule suspensions) involve elongated molecules that spontaneous align with boundaries as a result of steric interactions. Our analysis could be extended to allow for weak anchoring, with the introduction of an additional physical constant capturing the strength of anchoring Chandler and Spagnolie (2024b).
@@ II.4 Non-dimensionalization
$\bm{\sigma}^{h}=-p\,\mathbf{I}+2\,\mathbf{e},\qquad\bm{\sigma}^{a}=-S\mathbf{qq},\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\sigma}^{e}=-\xi^{2}\nabla\mathbf{q}\cdot\nabla\mathbf{q}^{\mathrm{T}},}$ (16)
@@
where $\lambda=\overline{\mu}/\mu$ is the viscosity ratio. The active stress coefficient in Eq. (16) is $S=\alpha/|\alpha|$, with $S=+1$ for an extensile nematic and $S=-1$ for a contractile nematic. The relative magnitude of elastic and active stresses is captured by the dimensionless parameter $\xi=\sqrt{K/|\alpha|}/R$, or ratio of the active length scale $\sqrt{K/|\alpha|}$ to the system size. The parameter $\xi$ is also related to the Ericksen number used in the literature on liquid crystals as $\mathrm{Er}=1/\xi^{2}$ Calderer and Liu (2000); Chandler and Spagnolie (2024a). Active stresses are negligible when $\xi\gg 1$, whereas they dominate when $\xi\to 0$, a regime where active turbulence is typically observed. In the present work, we assume that $\xi$ is of order unity.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{n}\cdot(\bm{\sigma}^{h}+\bm{\sigma}^{a}+\bm{\sigma}^{e}-\overline{\bm{\sigma}}^{h})+\mathbf{t}^{e}_{s}=\frac{1}{\mathrm{Ca}}\kappa\mathbf{n},}$ (18)
where $\mathbf{t}^{e}_{s}$ is given by Eq. (13) with $K$ replaced with $\xi^{2}$, and where we have introduced the active capillary number $\mathrm{Ca}=|\alpha|R/\gamma$ comparing the magnitudes of active versus capillary stresses. Under the assumption of vanishing Péclet number, the system is entirely governed by five parameters: the radius ratio $\eta$, the sign coefficient $S=\pm 1$ of the active stress, the capillary number $\mathrm{Ca}$, the dimensionless elastic constant $\xi$, and the viscosity ratio $\lambda$.
@@ III Linear stability analysis
We analyze the stability of small perturbations from the stable case of a circular drop centered at the origin. In the following, subscripts 0 and 1 are used to denote the base state and linear perturbations, respectively.
@@ III.1 Circular base state
$\beta_{0}=\theta+\frac{\pi}{2},$ (19)
which satisfies Eq. (15) along with the anchoring boundary conditions at $r=\eta$ and 111. The corresponding nematic director is $\smash{\mathbf{q}_{0}=\hat{\bm{\theta}}}$. Both fluids are still ($\mathbf{u}_{0}=\overline{\mathbf{u}}_{0}=\mathbf{0}$). The base-state active and elastic stress tensors in the outer nematic layer are given by
$\bm{\sigma}^{a}_{0}=-S\mathbf{q}_{0}\mathbf{q}_{0}=-S\hat{\bm{\theta}}\hat{\bm{\theta}},\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\sigma}^{e}_{0}=-\xi^{2}\nabla\mathbf{q}_{0}\cdot\nabla\mathbf{q}_{0}^{\mathrm{T}}=-\frac{\xi^{2}}{r^{2}}\hat{\bm{\theta}}\hat{\bm{\theta}}.}$ (20)
@@
$\overline{p}_{0}=S\ln\eta-\frac{\xi^{2}}{2\eta^{2}}+\frac{1}{\eta\mathrm{Ca}}.$ (23)